mathslover (mathslover):

sin ( 90 + theta ) = cos theta prove^

OpenStudy (anonymous):

i guess it depends on what you can use if you can use the addition angle formula, that would work if not maybe with a triangle you can see it

mathslover (mathslover):

i want to do with the triangle one but m wondering how will you show (90+theta) in a right angled triangle?

OpenStudy (anonymous):



mathslover (mathslover):

yes but that is 90 - theta? do 90-theta has any relation with 90 + theta ?

OpenStudy (anonymous):

\[\cos(x)=\frac{a}{c}=\sin(90-x)=-\sin(x-90)\] \[=\sin(x-90+180)=\sin(x+90)\]

mathslover (mathslover):

how : -sin(x-90) = sin(x-90+180) ?

OpenStudy (anonymous):

third equal sign because sine is odd, fourth because \(\sin(x)=-\sin(x+180)\)

mathslover (mathslover):

Gotcha now, thanks a lot!

OpenStudy (anonymous):

your on the other side of the unit circle 

OpenStudy (anonymous):

yw

mathslover (mathslover):

hey, how you did this step? : \[\large{-\sin(x-90) = sin(x-90+180) }\] ?

mathslover (mathslover):

@satellite73 did you subtract sin(180) both sides?

OpenStudy (anonymous):

i added 180 inside because \(\sin(x)=-\sin(x+180)\)

mathslover (mathslover):

oh k , got it!